OpenStudy (anonymous):

write in standard form: 2+3i all over 1+i what is the ans for this question

OpenStudy (anonymous):

\[\frac{2+3i}{1+i}=\frac{2+3i}{1+i}*\frac{1-i}{1-i}\] \[=\frac{2+3+i}{2}\] \[\frac{5}{2}+\frac{i}{2}\]

OpenStudy (anonymous):

thanks prophet